How many formula units are in 7.9 moles of lithium fluoride?
*Give your answer using scientific notation and no spaces. For example 6.02x10^23 or 8.96x10^24
*Round to the appropriate number of Significant Figures*

How many molecules are in 15 moles of carbon dioxide?
*Give your answer using scientific notation and no spaces. For example 6.02x10^23 or 8.96x10^24
*Round to the appropriate number of Significant Figures*

Convert 3.050x10^38 atoms of Iron to moles.
*Rounded to the HUNDREDTH*

Convert 1.20x10^24 formula units of NaCl to moles.
*Rounded to the HUNDREDTH*

Convert 6.02x10^22 molecules of Br2 to moles.
*Rounded to the HUNDREDTH*

Convert 1.00x10^9 formula units of PbO2 to moles.
*Rounded to the HUNDREDTH*
1
Expert's answer
2020-01-29T04:43:25-0500

N (LiCl) = 7.9 mol × 6.02 × 1023 = 47.558 × 1023 = 4.8 × 1024 of LiCl formula units

N(CO2) = 15 mol × 6.02 × 1023 = 90.3 × 1023 = 9.0 × 1024 of CO2 molecules

n (Fe) = 3.050 × 1038 / 6.02 × 1023 = 0.506644518 × 1015 = 5.07 × 1014 mol of Iron

n (NaCl) = 1.20 × 1024 / 6.02 × 1023 = 0.1993355 × 101 = 1.99 mol of NaCl formula units

n (Br2) = 6.02 × 1022 / 6.02 × 1023 = 1.00 × 10-1 mol of Br2 molecules

n (PbO2) = 1.00 × 109 / 6.02 × 1023 = 0.166113 × 10-14 = 1.66 × 10-15 mol of PbO2 formula units